# Thermometer: Towards Universal Calibration for Large Language Models

### 20 Feb 2024 v1

@@ Abstract
We consider the issue of calibration in large language models (LLM). Recent studies have found that common interventions such as instruction tuning often result in poorly calibrated LLMs. Although calibration is well-explored in traditional applications, calibrating LLMs is uniquely challenging. These challenges stem as much from the severe computational requirements of LLMs as from their versatility, which allows them to be applied to diverse tasks. Addressing these challenges, we propose thermometer, a calibration approach tailored to LLMs. thermometer learns an auxiliary model, given data from multiple tasks, for calibrating a LLM. It is computationally efficient, preserves the accuracy of the LLM, and produces better-calibrated responses for new tasks. Extensive empirical evaluations across various benchmarks demonstrate the effectiveness of the proposed method.
@@ 1 Introduction
Calibration is a desirable property of any probabilistic forecaster. Well-calibrated forecasts ensure that probabilities produced by the forecaster can be interpreted as accurate confidence estimates of the forecasts. Informally, this implies that the forecaster is more often wrong on predictions made with low probabilities than those made with high probabilities. Well-calibrated forecasts are crucial for both enabling trust in the forecaster’s predictions and incorporating the forecasts as part of a larger autonomous or semi-autonomous system.
@@
Large language models (LLMs) (Brown et al., 2020; Raffel et al., 2020; Touvron et al., 2023) define probability distributions over sequences of tokens and produce probabilistic forecasts over future tokens in a sequence and corresponding semantic entities represented by the tokens. Their ability to synthesize knowledge from large volumes of data, represented as sequences of tokens, has led to remarkable performance on diverse tasks, including question answering (Hendrycks et al., 2021), commonsense reasoning (Zhong et al., 2019), and machine translation (Zhu et al., 2020) among others. However, much like other probabilistic forecasters, before deploying LLMs in critical applications, it is important that they are well-calibrated in addition to being accurate. Unfortunately, a growing body of evidence suggests that while pre-trained LLMs are often well-calibrated, alignment interventions such as instruction tuning which make the pre-trained LLMs more usable, also harm calibration (OpenAI, 2023; Zhu et al., 2023).
@@
While calibration has long been studied (Brier, 1950; Dawid, 1982; Gneiting et al., 2007) and different approaches for improving calibration properties of probabilistic forecasters exist (Platt et al., 1999; Lakshminarayanan et al., 2017; Gal and Ghahramani, 2016; Guo et al., 2017), LLMs pose unique challenges. Training a LLM is expensive, and even inference typically incurs non-negligible expenses. This makes any calibration approach that requires multiple training runs prohibitively expensive. Even approaches that only require multiple inferences at test time can be unreasonably expensive for certain applications. Moreover, owing to their versatility, instruction-tuned LLMs are often applied, without further adaptation, to a diverse array of tasks. It is essential that methods for calibrating them do not affect the accuracy of the uncalibrated LLMs and that the calibration methods themselves can adapt to new tasks. Finally, measuring and alleviating calibration is challenging in cases where the LLM is required to generate free-form text. The equivalence class defined by the LLM-generated sequences that map to the same semantic content is large, making it challenging to assign meaningful confidence to the generation or even robustly assess the quality of the generation.
@@
We present thermometer for calibrating LLMs while alleviating the above challenges. thermometer learns, from multiple tasks, a parameterized mapping to map the outputs of the LLM to better-calibrated probabilities. Our approach is (i) computationally efficient: we do not require multiple training runs, and at inference time, we are only $\sim 0.5\%$ slower than the uncalibrated LLM, (ii) accuracy preserving: we build on temperature scaling (Guo et al., 2017) which ensures that the predictions do not change after our calibration procedure, and (iii) universal: once trained we do not require retraining when exposed to a new task. Moreover, similarly to recent work (Kadavath et al., 2022; Lin et al., 2022), we circumvent the challenges posed by free-form text generation by mapping the free-form text generation task to a next-token prediction task. We empirically evaluate thermometer on diverse benchmarks and models and find it consistently produce better-calibrated uncertainties than competing methods at a fraction of the computational cost. Moreover, we find thermometer transfers across datasets and model scales. thermometer trained for calibrating a smaller model (e.g., LLaMA-2-Chat 7B) also improves calibration of larger models (e.g., LLaMA-2-Chat 70B) from the same family of models.
@@ 2 Related Work
A variety of methods exist that aim to produce better-calibrated uncertainties. Post-hoc calibration methods learn to map the outputs of a pre-trained model to well-calibrated uncertainties. These include histogram binning (Zadrozny and Elkan, 2001; Naeini et al., 2015), isotonic regression (Zadrozny and Elkan, 2002), and approaches that assume parametric maps, including matrix, vector, and temperature scaling (Platt et al., 1999; Guo et al., 2017) as well as more ambitious variants (Kull et al., 2019) that aim to calibrate, in multi-class prediction problems, all class probabilities rather than just the probability of the most likely class (confidence calibration). The next-token prediction task is also a multi-class prediction problem, albeit one with a large number of classes. Here, the number of classes equals the number of tokens in the LLM’s vocabulary. This high dimensionality motivates us to focus on the more modest goal of confidence calibration.
@@
Other ab initio approaches involve training with label-smoothing (Szegedy et al., 2016), mix-up augmentations (zhang2017mixup), confidence penalties (Pereyra et al., 2017), focal loss (Mukhoti et al., 2020), or approximate Bayesian procedures (Izmailov et al., 2021). The substantial changes to the training process required by these approaches make them difficult to use with LLMs. Yet other approaches include ensembling over multiple models arrived at by retraining from different random initializations, for example, deep ensembles (Lakshminarayanan et al., 2017), or from perturbations of a single model, for example, Monte-Carlo dropout (Gal and Ghahramani, 2016). Training multiple LLM variants is prohibitively expensive, and while perturbations around a single model are possible, we find them to be not competitive with thermometer.
Jiang et al. (2021); Xiao et al. (2022); Chen et al. (2022) empirically evaluate calibration of LLMs, find evidence of miscalibration, and evaluate existing calibration interventions and combinations thereof with varying degrees of success. Park and Caragea (2022) use mixup while Desai and Durrett (2020) find temperature scaling and label smoothing effective for calibrating smaller encoder-only models. Others (Lin et al., 2022) employ supervised fine-tuning to produce verbalized uncertainties to be better calibrated on certain tasks. Such fine-tuning, however, is compute-intensive. An alternate body of work (Zhang et al., 2021; Kadavath et al., 2022; Mielke et al., 2022) approach calibration indirectly by learning an auxiliary model for predicting whether a generation is incorrect. Yet others (Tian et al., 2023), similar to our approach, consider RLHF-tuned LLMs and find that they can express better-calibrated uncertainties in the token space with carefully crafted prompts and that temperature scaling further improves calibration. Our work is orthogonal, extending temperature scaling to tasks without labeled data, and can be combined with their approach.
@@ 3 Background
using a large language model, ${\mathcal{M}}$, parameterized by weights, $\{{\mathcal{W}},{\bm{w}}_{1},\ldots,{\bm{w}}_{V}\}$. We use $p(y_{n}\mid{\bm{x}}_{n};{\mathcal{M}})$ as a notationally convenient stand in for $p(Y=y_{n}\mid X={\bm{x}}_{n};{\mathcal{M}})$, where $Y$ and $X$ are random variables corresponding to the completion and the prompt. We also note that $p(Y\mid X={\bm{x}}_{n};{\mathcal{M}})\in\Delta^{V}$ where $\Delta^{V}$ is a $V$-dimensional probability simplex. We view $\phi({\bm{x}}_{n},{\mathcal{W}})$ as a feature extractor that transforms the input prompt ${\bm{x}}_{n}$ into a $D$-dimensional representation. For notational brevity, we will suppress the dependence on ${\mathcal{W}}$ and use $\phi({\bm{x}}_{n},{\mathcal{W}})$ and $\phi({\bm{x}}_{n})$ interchangeably.
@@
$\text{acc}(B_{m})=1/|B_{m}|\sum_{n\in B_{m}}\mathbf{1}\left[y_{n}=\mathrm{argmax\;}p(Y|{\bm{x}}_{n};{\mathcal{M}})\right],$ and the average confidence within each bin,
$\text{conf}(B_{m})\!=\!1/|B_{m}|\sum_{n\in B_{m}}\!\!\mathrm{max\;}p(Y|{\bm{x}}_{n};\!{\mathcal{M}}\!),$
@@
Consider a multi-task setting. Given $K$ tasks with labeled datasets $\{{\mathcal{D}}_{k}\}_{k=1}^{K}=\{\{\left({\bm{x}}^{k}_{n},y^{k}_{n}\right)\}_{n=1}^{N_{k}}\}_{k=1}^{K}$, our goal is to learn to infer task-specific temperature parameters, in order to accurately infer the temperature for an unseen task given only $N_{*}$ prompts, ${\mathcal{D}}_{*}=\{{\bm{x}}_{n}^{*}\}_{n=1}^{N_{*}}$ but not the corresponding completions. To this end, we propose a probabilistic model with task-specific latent variables, $\tau_{k}$,
$\begin{split}p(\mathcal{D}_{1},\ldots\mathcal{D}_{K}\mid\nu_{0};{\mathcal{M}})=\prod_{k=1}^{K}\int p(\tau_{k}\mid\nu_{0})\prod_{n=1}^{N_{k}}p(y_{n}^{k}\mid{\bm{x}}_{n}^{k},\tau_{k};{\mathcal{M}})d\tau_{k},\end{split}$ (1)
where $p(y_{n}^{k}\mid{\bm{x}}_{n}^{k},\tau_{k};{\mathcal{M}})=\frac{\text{exp}({\bm{w}}_{y_{n}^{k}}^{T}\phi({\bm{x}}_{n}^{k};{\mathcal{W}})/\tau_{k})}{\sum_{v^{\prime}=1}^{V}\text{exp}({\bm{w}}_{v^{\prime}}^{T}\phi({\bm{x}}_{n}^{k};{\mathcal{W}})/\tau_{k})}$, and $p(\tau_{k}\mid\nu_{0})$ is a prior on $\tau_{k}$ for regularizing it away from zero and large positive values.
@@ 4.1 A Variational Lower-bound
Figure 1: Evidence accumulation. We illustrate evidence accumulation for the Professional Law dataset from MMLU. It contains $N_{k}=1533$ instances. The left panel, shows twenty of the possible $1533$ Gaussians, ${\mathcal{N}}(\tau_{k}\mid\psi_{\bm{\theta}}(\phi({\bm{x}}_{n}^{k})),\epsilon=0.01)$. The right panel, plotw the Gaussian proportional to $\prod_{n=1}^{1533}{\mathcal{N}}(\tau_{k}\mid\psi_{\bm{\theta}}(\phi({\bm{x}}_{n}^{k})),\epsilon=0.01)$. The resulting Gaussian is nearly a point mass at $\frac{1}{1533}\sum_{n=1}^{1533}\psi_{\bm{\theta}}(\phi({\bm{x}}_{n}^{k}))$.
@@
$\begin{split}{\mathcal{L}}(\bm{\theta})=\sum_{k=1}^{K}\mathbb{E}_{q(\tau_{k};\bm{\theta})}[\log p\left(\mathcal{D}_{k}\mid\tau_{k};{\mathcal{M}}\right)]-\mathrm{KL}\left(q(\tau_{k};\bm{\theta})\;||\;p(\tau_{k}\mid\nu_{0})\right),\end{split}$ (2)
where $q(\tau_{k};\bm{\theta})$ is a variational approximation to the posterior distribution $p(\tau_{k}\mid\mathcal{D}_{1},\ldots,\mathcal{D}_{k})$. We choose a product of Gaussians form for the variational approximation. Each Gaussian in the product has a fixed variance $\epsilon$, and a mean parameterized by a recognition network (for example, a MLP), $\psi_{\bm{\theta}}:^{D}\rightarrow_{+}$, where $\bm{\theta}\in^{P}$ represents the parameters of the recognition network that are shared across $\mathcal{D}_{1},\ldots,\mathcal{D}_{k}$. The resulting variational approximation is,
$\begin{split}q(\tau_{k};\bm{\theta})\propto\prod_{n=1}^{N_{k}}{\mathcal{N}}(\tau_{k}\mid\psi_{\bm{\theta}}(\phi({\bm{x}}_{n}^{k})),\epsilon)={\mathcal{N}}\bigg{(}\tau_{k}\mid\frac{1}{N_{k}}\sum_{n=1}^{N_{k}}\psi_{\bm{\theta}}(\phi({\bm{x}}_{n}^{k})),\frac{\epsilon}{N_{k}}\bigg{)},\end{split}$ (3)
@@
$\begin{split}{\mathcal{L}}(\bm{\theta})=\sum_{k=1}^{K}\mathbb{E}_{q(\tau_{k};\bm{\theta})}[\log p\left(\mathcal{D}_{k}|\tau_{k};{\mathcal{M}}\right)+\log p(\tau_{k}|\nu_{0})]+\text{C},\end{split}$ (4)
@@
~$\begin{split}\tilde{{\mathcal{L}}}(\bm{\theta})=\sum_{k=1}^{K}\log p\left(\mathcal{D}_{k}\;\bigg{|}\;\frac{1}{N_{k}}\sum_{n=1}^{N_{k}}\psi_{\bm{\theta}}(\phi({\bm{x}}_{n}^{k}));{\mathcal{M}}\right)-\lambda_{\text{reg}}\log p\left(\frac{1}{N_{k}}\sum_{n=1}^{N_{k}}\psi_{\bm{\theta}}(\phi({\bm{x}}_{n}^{k}))\;\bigg{|}\;\nu_{0}\right),\end{split}$ (5)
where we have dropped the constant, plugged in the point mass approximation in Equation 4, and introduced a hyperparameter $\lambda_{\text{reg}}$ for balancing the prior and the likelihood terms. We place a Gamma prior, with a mode at one, on the temperatures $\tau_{k}$, $p(\tau_{k}\mid\nu_{0})=\text{Gamma}(\tau_{k}\mid{\alpha_{0}=1.25,\beta_{0}=4})$ [^2] We use the shape scale parameterization, with shape = 1.25 and scale = 4.. This prior enforces non-negative temperatures, and although the prior’s mode at one encodes our belief that ${\mathcal{M}}$ is already reasonably calibrated, the high prior variance emphasizes that this is only a weakly held belief. Finally, we train thermometer by minimizing, $\bm{\theta}^{*}=\underset{\bm{\theta}\in^{P}}{\operatorname*{arg\,min}}-\tilde{{\mathcal{L}}}(\bm{\theta})$, or equivalently maximizing an approximate333the approximation is increasingly accurate with larger $N_{k}$. lower bound to the log-marginal likelihood. We use stochastic optimization with mini-batched gradients to optimize ~$\tilde{{\mathcal{L}}}(\bm{\theta})$. Algorithm 1 summarizes our learning procedure. The temperature estimation step in the algorithm involves averaging over the minibatch which introduces a small bias in the minibatch gradients. However, this bias decreases linearly with increasing batch size (Lemma 4.1), and disappears in the full-batch limit. Empirically, we find performance to be largely invariant to batch size (Table 8).
@@
$\triangleright$ Initialize thermometer, $\bm{\theta}_{0}\leftarrow\bm{\theta}_{\text{init}}$; Denote the number of minibatches by $B$.

### 27 Jun 2024 v2

@@ Abstract
We consider the issue of calibration in large language models (LLM). Recent studies have found that common interventions such as instruction tuning often result in poorly calibrated LLMs. Although calibration is well-explored in traditional applications, calibrating LLMs is uniquely challenging. These challenges stem as much from the severe computational requirements of LLMs as from their versatility, which allows them to be applied to diverse tasks. Addressing these challenges, we propose thermometer, a calibration approach tailored to LLMs. thermometer learns an auxiliary model, given data from multiple tasks, for calibrating a LLM. It is computationally efficient, preserves the accuracy of the LLM, and produces better-calibrated responses for new tasks. Extensive empirical evaluations across various benchmarks demonstrate the effectiveness of the proposed method [^1] The code is available at https://github.com/maohaos2/Thermometer..
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Calibration performance against inference runtime. Different methods for calibrating LLaMA-2-Chat 7B compared on the Professional Law task of MMLU on a A100, 40 GB GPU. The task contains $1533$ questions. Our method, thermometer is significantly faster than methods that require multiple forward passes (Wei & Zou, 2019; Xiong et al., 2023; Jiang et al., 2023a) at inference time and achieves lower calibration error compared to methods with comparable runtime (Tian et al., 2023). Vanilla refers to the no-calibration baseline, and TS-CV is a temperature scaling variant (Section 5). Similar trends hold for other benchmarks.
@@
Well-calibrated forecasts are a desirable property of any probabilistic forecaster. They ensure that probabilities produced by the forecaster can be interpreted as accurate confidence estimates of the forecasts. Informally, this implies that the forecaster is more often wrong on predictions made with low probabilities than those made with high probabilities. Such forecasts are useful for both enabling trust in the forecaster’s predictions and incorporating the forecasts as part of a larger autonomous or semi-autonomous system.
@@
Large language models (LLMs) (Brown et al., 2020; Raffel et al., 2020; Touvron et al., 2023) define probability distributions over sequences of tokens and produce probabilistic forecasts over future tokens in a sequence222and corresponding semantic entities represented by the tokens.. Their ability to synthesize knowledge from large volumes of data, represented as sequences of tokens, has led to remarkable performance on diverse tasks, including question answering (Hendrycks et al., 2020), commonsense reasoning (Zhong et al., 2019), and machine translation (Zhu et al., 2020) among others. However, much like other probabilistic forecasters, before deploying LLMs in critical applications, it is important that they are well-calibrated in addition to being accurate. Unfortunately, a growing body of evidence suggests that while pre-trained LLMs are often well-calibrated, alignment interventions such as instruction tuning which make the pre-trained LLMs more usable, also harm calibration (OpenAI, 2023; Zhu et al., 2023).
@@
While calibration has long been studied (Brier, 1950; Dawid, 1982; Gneiting et al., 2007) and different approaches for improving calibration properties of probabilistic forecasters exist (Platt et al., 1999; Lakshminarayanan et al., 2017; Gal & Ghahramani, 2016; Guo et al., 2017), LLMs pose unique challenges. Training a LLM is expensive, and even inference typically incurs non-negligible expenses. This makes any calibration approach that requires multiple training runs prohibitively expensive. Even approaches that only require multiple inferences at test time can be unreasonably expensive for certain applications. Moreover, owing to their versatility, instruction-tuned LLMs are often applied, without further adaptation, to a diverse array of tasks. It is essential that methods for calibrating them do not affect the accuracy of the uncalibrated LLMs and that the calibration methods themselves can adapt to new tasks. Finally, measuring and alleviating calibration is challenging in cases where the LLM is required to generate free-form text. The equivalence class defined by the LLM-generated sequences that map to the same semantic content is large, making it challenging to assign meaningful confidence to the generation or even robustly assess the quality of the generation.
@@
We present thermometer for calibrating LLMs while alleviating the above challenges. thermometer learns, from multiple tasks, a parameterized mapping to map the outputs of the LLM to better-calibrated probabilities. Our approach is (i) computationally efficient: we do not require multiple training runs, and at inference time, we are only $\sim 0.5\%$ slower than the uncalibrated LLM, (ii) accuracy preserving: we build on temperature scaling (Guo et al., 2017) which provably guarantees that greedy-decoded predictions do not change after our calibration procedure, and (iii) takes a step towards being universal: once trained we do not require retraining when exposed to a similar but new task. Moreover, similarly to recent work (Kadavath et al., 2022; Lin et al., 2022), we circumvent the challenges posed by free-form text generation by mapping the free-form text generation task to a next-token prediction task. We empirically evaluate thermometer on diverse benchmarks and models and find it consistently produce better-calibrated uncertainties than competing methods at a fraction of the computational cost. Moreover, we find thermometer transfers across datasets and model scales. thermometer trained for calibrating a smaller model (e.g., LLaMA-2-Chat 7B) also improves calibration of larger models (e.g., LLaMA-2-Chat 70B) from the same family of models.
@@ 2 Related Work
A variety of methods exist that aim to produce better-calibrated uncertainties. Post-hoc calibration methods learn to map the outputs of a pre-trained model to well-calibrated uncertainties. These include histogram binning (Zadrozny & Elkan, 2001; Naeini et al., 2015), isotonic regression (Zadrozny & Elkan, 2002), and approaches that assume parametric maps, including matrix, vector, and temperature scaling (Platt et al., 1999; Guo et al., 2017) as well as more ambitious variants (Kull et al., 2019) that aim to calibrate, in multi-class prediction problems, all class probabilities rather than just the probability of the most likely class (confidence calibration). The next-token prediction task is also a multi-class prediction problem, albeit one with a large number of classes. Here, the number of classes equals the number of tokens in the LLM’s vocabulary. This high dimensionality motivates us to focus on the more modest goal of confidence calibration.
@@
Other ab initio approaches involve training with label-smoothing (Szegedy et al., 2016), mix-up augmentations (Zhang et al., 2017), confidence penalties (Pereyra et al., 2017), focal loss (Mukhoti et al., 2020), or approximate Bayesian procedures (Izmailov et al., 2021). The substantial changes to the training process required by these approaches make them difficult to use with LLMs. Yet other approaches include ensembling over multiple models arrived at by retraining from different random initializations, for example, deep ensembles (Lakshminarayanan et al., 2017), or from perturbations of a single model, for example, Monte-Carlo dropout (Gal & Ghahramani, 2016). Training multiple LLM variants is prohibitively expensive, and while perturbations around a single model are possible, we find them to be not competitive with thermometer.
Jiang et al. (2021); Xiao et al. (2022); Chen et al. (2022) empirically evaluate calibration of LLMs, find evidence of miscalibration, and evaluate existing calibration interventions and combinations thereof with varying degrees of success. Park & Caragea (2022) use mixup while Desai & Durrett (2020) find temperature scaling and label smoothing effective for calibrating smaller encoder-only models. Others (Lin et al., 2022) employ supervised fine-tuning to produce verbalized uncertainties to be better calibrated on certain tasks. Such fine-tuning, however, is compute-intensive. An alternate body of work (Zhang et al., 2021; Kadavath et al., 2022; Mielke et al., 2022) approach calibration indirectly by learning an auxiliary model for predicting whether a generation is incorrect. There is also a body of work (Abbas et al., 2024; Han et al., 2023; Jiang et al., 2023b; Zhao et al., 2021; Zhou et al., 2024) that use the term calibration to mean de-biasing the predictions of a language model to biases introduced by the choice and ordering of in-context examples. This is distinct from our notion of statistical calibration. Yet others (Tian et al., 2023; Xiong et al., 2023), similar to our approach, consider RLHF-tuned LLMs and find that they can express better-calibrated uncertainties with carefully crafted prompts. Our work is orthogonal, extending temperature scaling to tasks without labeled data, and can be combined with their approach.
@@ 3 Background
using a large language model, ${\mathcal{M}}$, parameterized by weights, $\{{\mathcal{W}},{\bm{w}}_{1},\ldots,{\bm{w}}_{V}\}$. We use $p(y_{n}\mid{\bm{x}}_{n};{\mathcal{M}})$ as a notationally convenient stand in for ${p(Y=y_{n}|X={\bm{x}}_{n};{\mathcal{M}})}$, where $Y$ and $X$ are random variables corresponding to the completion and the prompt. We also note that ${p(Y\mid X={\bm{x}}_{n};{\mathcal{M}})\in\Delta^{V}}$ where $\Delta^{V}$ is a $V$-dimensional probability simplex. We view $\phi({\bm{x}}_{n},{\mathcal{W}})$ as a feature extractor that transforms the input prompt ${\bm{x}}_{n}$ into a $D$-dimensional representation. For notational brevity, we will suppress the dependence on ${\mathcal{W}}$ and use $\phi({\bm{x}}_{n},{\mathcal{W}})$ and $\phi({\bm{x}}_{n})$ interchangeably.
@@
$\small\text{acc}(B_{m})=1/|B_{m}|\sum_{n\in B_{m}}\mathbf{1}\left[y_{n}=\mathrm{argmax\;}p(Y|{\bm{x}}_{n};{\mathcal{M}})\right],$ and the average confidence within each bin,
$\small\text{conf}(B_{m})\!=\!1/|B_{m}|\sum_{n\in B_{m}}\!\!\mathrm{max\;}p(Y|{\bm{x}}_{n};\!{\mathcal{M}}\!),$
@@
Consider a multi-task setting. Given $K$ tasks with labeled datasets $\{{\mathcal{D}}_{k}\}_{k=1}^{K}=\{\{\left({\bm{x}}^{k}_{n},y^{k}_{n}\right)\}_{n=1}^{N_{k}}\}_{k=1}^{K}$, our goal is to learn to infer task-specific temperature parameters, in order to accurately infer the temperature for an unseen task given only $N_{*}$ prompts, ${\mathcal{D}}_{*}=\{{\bm{x}}_{n}^{*}\}_{n=1}^{N_{*}}$ but not the corresponding completions. To this end, we propose a probabilistic model with task-specific latent variables, $\tau_{k}$, for modeling the conditional distribution of completions given a prompt,
$\begin{split}p(&\mathcal{D}_{1},\ldots\mathcal{D}_{K}\mid\nu_{0};{\mathcal{M}})\\ &=\prod_{k=1}^{K}\int p(\tau_{k}\mid\nu_{0})\prod_{n=1}^{N_{k}}p(y_{n}^{k}\mid{\bm{x}}_{n}^{k},\tau_{k};{\mathcal{M}})d\tau_{k},\end{split}$ (1)
where $p(y_{n}^{k}\mid{\bm{x}}_{n}^{k},\tau_{k};{\mathcal{M}})=\frac{\text{exp}({\bm{w}}_{y_{n}^{k}}^{T}\phi({\bm{x}}_{n}^{k};{\mathcal{W}})/\tau_{k})}{\sum_{v^{\prime}=1}^{V}\text{exp}({\bm{w}}_{v^{\prime}}^{T}\phi({\bm{x}}_{n}^{k};{\mathcal{W}})/\tau_{k})}$, $p(\tau_{k}\mid\nu_{0})$ is a prior on $\tau_{k}$ for regularizing it away from zero and large positive values, and $\nu_{0}$ is the prior’s hyper-parameters.
@@ 4.1 A Variational Lower-bound
Figure 2: Evidence accumulation. We illustrate evidence accumulation for the Professional Law dataset from MMLU. It contains $N_{k}=1533$ instances. The left panel, shows twenty of the possible $1533$ Gaussians, ${\mathcal{N}}(\tau_{k}\mid\psi_{\bm{\theta}}(\phi({\bm{x}}_{n}^{k})),\epsilon=0.01)$. The right panel, plots the Gaussian proportional to $\prod_{n=1}^{1533}{\mathcal{N}}(\tau_{k}\mid\psi_{\bm{\theta}}(\phi({\bm{x}}_{n}^{k})),\epsilon=0.01)$. The resulting Gaussian is nearly a point mass at $\frac{1}{1533}\sum_{n=1}^{1533}\psi_{\bm{\theta}}(\phi({\bm{x}}_{n}^{k}))$.
@@
$\begin{split}{\mathcal{L}}(\bm{\theta})=\sum_{k=1}^{K}&\mathbb{E}_{q(\tau_{k};\bm{\theta})}[\log p\left(\mathcal{D}_{k}\mid\tau_{k};{\mathcal{M}}\right)]\\ &-\mathrm{KL}\left(q(\tau_{k};\bm{\theta})\;||\;p(\tau_{k}\mid\nu_{0})\right),\end{split}$ (2)
where $q(\tau_{k};\bm{\theta})$ is a variational approximation to the posterior distribution $p(\tau_{k}\mid\mathcal{D}_{1},\ldots,\mathcal{D}_{k})$. We choose a product of Gaussian PDF for the variational approximation. Each corresponding Gaussian variable has a fixed variance $\epsilon$, and a mean parameterized by a recognition network (for example, a MLP), $\psi_{\bm{\theta}}:^{D}\rightarrow_{+}$, where $\bm{\theta}\in^{P}$ represents the parameters of the recognition network that are shared across $\mathcal{D}_{1},\ldots,\mathcal{D}_{k}$. The resulting variational approximation is,
$\begin{split}q(\tau_{k};\bm{\theta})&\propto\prod_{n=1}^{N_{k}}{\mathcal{N}}(\tau_{k}\mid\psi_{\bm{\theta}}(\phi({\bm{x}}_{n}^{k})),\epsilon),\\ &{\mathcal{N}}\bigg{(}\tau_{k}\large\mid\frac{1}{N_{k}}\sum_{n=1}^{N_{k}}\psi_{\bm{\theta}}(\phi({\bm{x}}_{n}^{k})),\frac{\epsilon}{N_{k}}\bigg{)},\end{split}$ (3)
@@
${\mathcal{L}}(\bm{\theta})=\sum_{k=1}^{K}\mathbb{E}_{q(\tau_{k};\bm{\theta})}[\log p\left(\mathcal{D}_{k}|\tau_{k};{\mathcal{M}}\right)+\log p(\tau_{k}|\nu_{0})]+\text{C},$ (4)
@@
~$\begin{split}\tilde{{\mathcal{L}}}(\bm{\theta})=\sum_{k=1}^{K}&\log p\left(\mathcal{D}_{k}\;\bigg{|}\;\frac{1}{N_{k}}\sum_{n=1}^{N_{k}}\psi_{\bm{\theta}}(\phi({\bm{x}}_{n}^{k}));{\mathcal{M}}\right)\\ -&\lambda_{\text{reg}}\log p\left(\frac{1}{N_{k}}\sum_{n=1}^{N_{k}}\psi_{\bm{\theta}}(\phi({\bm{x}}_{n}^{k}))\;\bigg{|}\;\nu_{0}\right),\end{split}$ (5)
where we have dropped the constant, plugged in the point mass approximation in Equation 4, and introduced a hyperparameter $\lambda_{\text{reg}}$ for balancing the prior and the likelihood terms. We place a Gamma prior, with a mode at one, on the temperatures $\tau_{k}$, $p(\tau_{k}\mid\nu_{0})=\text{Gamma}(\tau_{k}\mid{\alpha_{0}=1.25,\beta_{0}=4})$ [^4] We use the shape scale parameterization, with shape = 1.25 and scale = 4.. This prior enforces non-negative temperatures, and although the prior’s mode at one encodes our belief that ${\mathcal{M}}$ is already reasonably calibrated, the high prior variance emphasizes that this is only a weakly held belief. Finally, we train thermometer by minimizing, $\bm{\theta}^{*}=\underset{\bm{\theta}\in^{P}}{\operatorname*{arg\,min}}-\tilde{{\mathcal{L}}}(\bm{\theta})$, or equivalently maximizing an approximate555the approximation is increasingly accurate with larger $N_{k}$. lower bound to the log-marginal likelihood. We use stochastic optimization with mini-batched gradients to optimize ~$\tilde{{\mathcal{L}}}(\bm{\theta})$. Algorithm 1 summarizes our learning procedure. The temperature estimation step in the algorithm involves averaging over the minibatch which introduces a small bias in the minibatch gradients. However, this bias decreases linearly with increasing batch size (Lemma 4.1), and disappears in the full-batch limit. Empirically, we find performance to be largely invariant to batch size (Table 8).
@@
$\triangleright$ Initialize thermometer, $\bm{\theta}_{0}\leftarrow\bm{\theta}_{\text{init}}$; Denote the number of mini-batches by $B=\lceil\frac{N}{N_{b}}\rceil$, where $N=|\mathrm{D}^{\mathrm{dev}}|$.
@@
       $\triangleright$ Sample, uniformly at random, a task $k$ and its dataset ${\mathcal{D}}^{dev}_{k}$. $\triangleright$ Sample a batch of $N_{b}$ samples from ${\mathcal{D}}^{dev}_{k}$, i.e., $\{\left({\bm{x}}_{n}^{k},y_{n}^{k}\right)\}_{n=1}^{N_{b}}$. $\triangleright$ Estimate temperature $\hat{\tau}_{k}=\frac{1}{N_{b}}\sum_{n=1}^{N_{b}}\psi_{\bm{\theta}}(\phi({\bm{x}}_{n}^{k}))$. $\triangleright$ Evaluate the loss w.r.t batch of samples, i.e., ~$-\tilde{{\mathcal{L}}}(\bm{\theta})_{N_{b}}=-\sum_{n=1}^{N_{b}}\log p(y_{n}^{k}\mid{\bm{x}}_{n}^{k},\hat{\tau}_{k};{\mathcal{M}})-\frac{\lambda_{\text{reg}}}{B}\cdot\log p(\hat{\tau}_{k}\mid\nu_{0})$triangleright Update the thermometer parameter using AdamW, i.e., $\bm{\theta}_{m}\leftarrow\text{AdamW}(\bm{\theta}_{m-1},\gamma,-\nabla_{\bm{\theta}}\tilde{{\mathcal{L}}}(\bm{\theta})_{b})$. if $m>\mathrm{burnin}$ then              $\triangleright$ Checkpoint $\bm{\theta}_{m}$ every $m^{\prime}$ iterations.
       end if
end for $\triangleright$ Evaluate ~$\tilde{{\mathcal{L}}}(\bm{\theta})$ on $\mathrm{D}^{\mathrm{val}}$ for each check-pointed $\bm{\theta}$ and set $\bm{\theta}^{\ast}$ to the check-pointed $\bm{\theta}$ with lowest ~$-\tilde{{\mathcal{L}}}(\bm{\theta})$.
output Optimized thermometer, $\bm{\theta}^{\ast}.$